OpenStudy (anonymous):

Divide: 3y + 12/8y^3 divide 9y + 36/16y^3

OpenStudy (anonymous):

16y^3 is divided by 36 only or whole?

OpenStudy (anonymous):

36 over 16y^3

OpenStudy (anonymous):

(9y+36/16y^3)/(3y+12/8y^3) =[(144y^4+36)/16y^3] / [(24y^4+12)/8y^3] =[36(4y^4+1)/16y^3] / [12(2y^4+1)/8y^3] =[3(4y^4+1)] / (2y^4+1) =12y^4+3/2y^4+1 =6-3/2y^4+1

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes